OpenStudy (anonymous):

how do i evaluate f(x) = log(8)x for 64?

OpenStudy (a_clan):

f(64) = log(8) 64

OpenStudy (anonymous):

I'm confused...the question is asking for a y value with 64 as x value...

OpenStudy (anonymous):

i need to know what the y value is I can do it once someone explains how

OpenStudy (a_clan):

y = f(x)

OpenStudy (a_clan):

\[Y_{64} = f(64) = \log(8) 64\]

OpenStudy (a_clan):

=log(8) (8^2) =2 log(8) 8 =2

OpenStudy (anonymous):

ohhh okay now I get it and I feel stupid for not getting it at first xD Thank you!

OpenStudy (a_clan):

np